OpenStudy (anonymous):

Medal to the first+ right answer :D Solve for y. 3y – 12 = –6 A. 2 B. 18 C. –54 D. –6

OpenStudy (anonymous):

2 is the correct answeer

OpenStudy (anonymous):

thank you so much Kinnybowlin :)

OpenStudy (anonymous):

Add 12 to each side: \[3y - 12 + 12 = -6 + 12\] \[3y = 6\] Divide each side by 3: \[\frac{ 3y }{ 3 } = \frac{ 6 }{ 3 }\] \[y = 2\]

OpenStudy (anonymous):

Kimmy*****lol i just got my glasses off my e i didnd read it right

OpenStudy (anonymous):

@kimmybowlin It is against the OpenStudy CoC (Code of Conduct) to directly give the answer. You are supposed to guide the asker and not just give the answer.

OpenStudy (anonymous):

wat did u not read right??? because u gave the explanation i stop on mine i know wat to do